# Packing measure and dimension of the limit sets of IFSs of generalized complex continued fractions1112020 Mathematics Subject Classification: 28A78, 28A80

### v5

@@ Abstract
We consider a family of conformal iterated function systems (for short, CIFSs) of generalized complex continued fractions. Note that in our previous paper we showed that the proper-dimensional Hausdorff measure of the limit set is zero and the packing measure of the limit set with respect to the Hausdorff dimension is positive. In this paper, we show that the packing dimension and the Hausdorff dimension of the limit set of each CIFS in the family are equal, and the proper-dimensional packing measure of the limit set is finite.  [^3] Keywords: infinite conformal iterated function systems, fractal geometry, packing measures, packing dimension, generalized complex continued fractions.
@@ 1 Introduction and the main results
Mathematically speaking, iterated function systems are powerful tools to construct fractals (more precisely, limit sets) and there are many papers on the study of estimating the dimensions and measures of the limit sets constructed by iterated function systems. For example, Falconer’s paper [1] and book [2] consider the general theory and examples of the limit sets constructed by conformal iterated function systems with finitely many mappings (for short, finite CIFSs) on a interval under the open set condition (for short, the OSC), and discuss the dimensions and measures of the limit sets. In addition, Mauldin’s and Urbański’s paper [7] presents the general theory of estimating the dimensions and measures of the limit sets constructed by finite CIFSs on general compact connected subset of $\mathbb{R}^{d}$ under the OSC (and some conditions). By the formula and theorems in [1], [2] and [7], under the OSC (and some conditions), we deduce that the packing dimension of the limit set constructed by any finite CIFS equals the Hausdorff dimension of the limit set, the proper-dimensional Hausdorff measure of the limit set is positive and finite, and so is the proper-dimensional packing measure of the limit set. Here, it should be mentioned that the positivity and finiteness of the Hausdorff measure of the limit set of any finite CIFS with respect to the zero of the pressure function is equivalent to the OSC (see [12]), and there is an example of a self-similar set with zero Hausdorff measure but positive and finite packing measure without the OSC (see [13]).
In addition, Mauldin’s and Urbański’s paper [7] presents the general theory of estimating the dimensions and measures of the fractals (more precisely, the limit sets too) constructed by conformal iterated function systems with infinitely many mappings (for short, infinite CIFSs) under the OSC (and some conditions). Indeed, they generalize the above formula and theorems, and by the generalized theorems in their paper we may obtain non-positiveness of the proper-dimensional Hausdorff measure of the limit sets constructed by infinite CIFSs even if we assume the OSC. This theorem indicates that we may find a different phenomenon of infinite CIFSs from finite CIFSs under the OSC. Here, it should be mentioned that there are now many directions for the studies of the limit sets constructed by infinite CIFSs with the OSC (for example, see [8], [5], [9], [14] and so on) or without the OSC (for example, see [10], [11], and so on).
@@
Let $\hat{J}$ be the limit set of $\hat{S}$ (see Definition 2.1) and $\hat{h}$ be the Hausdorff dimension of $\hat{J}$. For each $s\geq 0$, we denote by $\mathcal{H}^{s}$ the $s$-dimensional Hausdorff measure and denoted by $\mathcal{P}^{s}$ the $s$-dimensional packing measure. Regarding this example, Mauldin and Urbański showed the following theorem.
@@
Note that they obtained an example of infinite CIFS for which the proper-dimensional Hausdorff measure of the limit set is zero and the proper-dimensional packing measure of the limit set is positive and finite. That is, they found an example of a phenomenon of infinite CIFSs with OSC which cannot hold in any finite CIFSs with the OSC.
In our previous papers [3] and [4], we considered a family of CIFSs of generalized complex continued fractions in order to present another examples of infinite CIFSs. In particular, we estimated the Hausdorff dimension of the limit set of each CIFS of the generalized complex continued fractions, and showed non-positiveness of the Hausdorff measure and positiveness of the packing measure of the limit set with respect to the Hausdorff dimension of the limit set. Note that the family of the CIFSs introduced in the papers [3] and [4] has uncountably many elements. However, there are unsolved problems in [3] and [4]: the estimation of the packing dimension and the proper-dimensional packing measure of the limit set. Therefore, the aim of this paper is to estimate the packing dimension and the proper-dimensional packing measure of each limit set in the family of the CIFSs of the generalized complex continued fractions, in order to improve the previous works [3] and [4].
@@
By the general theory of finite CIFSs with the OSC, the proper-dimensional Hausdorff measure and the proper-dimensional packing measure of the limit set is positive and finite (see [1], [2] and [7]). However, Corollary 1.5 indicates that for each $S_{\tau}$ in the family of CIFSs of generalized complex continued fractions which consists of uncountably many elements, the packing dimension of the limit set equals the Hausdorff dimension of the limit set, the proper-dimensional Hausdorff measure of the limit set is zero, and the proper-dimensional packing measure of the limit set is positive and finite. This is the phenomenon which Mauldin and Urbański first discovered in [7] and cannot hold in the finite CIFSs under the OSC.
@@
$m_{\tau}(B(x,r))$ $\geq m_{\tau}(B(0,\tilde{r}))\geq m_{\tau}\left(\bigcup_{a\in I_{\tau}(\tilde{r})}\phi_{a}(X)\right)\gtrsim\sum_{a\in I_{\tau}(\tilde{r})}\ |a|^{-2h_{\tau}}$
#$\gtrsim\#(I_{\tau}(\tilde{r}))\cdot r^{2h_{\tau}}\gtrsim r^{2h_{\tau}-2}>r^{h_{\tau}},$
@@ 3 CIFSs of generalized complex continued fractions
$|\phi_{a}(z)|\cdot|a|=\frac{|a|}{|a+z|}=\sqrt{\frac{|a|^{2}}{|a+z|^{2}}}=\sqrt{\frac{|\phi_{a}^{\prime}(z)|}{|\phi_{a}^{\prime}(0)|}}\leq\sqrt{C}\leq K_{0}\leq K.$

### v6

@@ Abstract
We consider a family of conformal iterated function systems (for short, CIFSs) of generalized complex continued fractions. Note that in our previous paper we showed that the proper-dimensional Hausdorff measure of the limit set of each CIFS is zero and the packing measure of the limit set with respect to the Hausdorff dimension is positive. In this paper, we show that the packing dimension and the Hausdorff dimension of the limit set of each CIFS in the family are equal, and the proper-dimensional packing measure of the limit set is finite.  [^3] Keywords: infinite conformal iterated function systems, fractal geometry, packing measures, packing dimension, generalized complex continued fractions.
@@ 1 Introduction and the main results
Mathematically speaking, iterated function systems are powerful tools to construct fractals (more precisely, limit sets) and there are many papers on the study of estimating the dimensions and measures of the limit sets constructed by iterated function systems. For example, Falconer’s paper [1] and book [2] consider the general theory and examples of the limit sets constructed by conformal iterated function systems with finitely many mappings (for short, finite CIFSs) on intervals under the open set condition (for short, the OSC), and discuss the dimensions and measures of the limit sets. In addition, Mauldin’s and Urbański’s paper [7] presents the general theory of estimating the dimensions and measures of the limit sets constructed by finite CIFSs on general compact connected subsets of $\mathbb{R}^{d}$ under the OSC (and some conditions). By the formula and theorems in [1], [2] and [7], under the OSC (and some conditions), we deduce that the packing dimension of the limit set constructed by any finite CIFS equals the Hausdorff dimension of the limit set, the proper-dimensional Hausdorff measure of the limit set is positive and finite, and so is the proper-dimensional packing measure of the limit set. Here, it should be mentioned that the positivity and finiteness of the Hausdorff measure of the limit set of any finite CIFS with respect to the zero of the pressure function are equivalent to the OSC (see [12]), and there is an example of a self-similar set with zero Hausdorff measure but positive and finite packing measure without the OSC (see [13]).
In addition, Mauldin’s and Urbański’s paper [7] presents the general theory of estimating the dimensions and measures of the fractals (more precisely, the limit sets too) constructed by conformal iterated function systems with infinitely many mappings (for short, infinite CIFSs) under the OSC (and some conditions). Indeed, they generalize the above formula and theorems, and by the generalized theorems in their paper we may obtain the non-positivity of the proper-dimensional Hausdorff measure of the limit sets constructed by infinite CIFSs even if we assume the OSC. This theorem indicates that we may find a different phenomenon of infinite CIFSs from finite CIFSs under the OSC. Here, it should be mentioned that there are now many directions for the studies of the limit sets constructed by infinite CIFSs with the OSC (for example, see [8], [5], [9], [14] and so on) or without the OSC (for example, see [10], [11], and so on).
@@
Let $\hat{J}$ be the limit set of $\hat{S}$ (see Definition 2.1) and $\hat{h}$ be the Hausdorff dimension of $\hat{J}$. For each $s\geq 0$, we denote by $\mathcal{H}^{s}$ the $s$-dimensional Hausdorff measure and denote by $\mathcal{P}^{s}$ the $s$-dimensional packing measure. Regarding this example, Mauldin and Urbański showed the following theorem.
@@
Note that they obtained an example of infinite CIFSs for which the proper-dimensional Hausdorff measure of the limit set is zero and the proper-dimensional packing measure of the limit set is positive and finite. That is, they found an example of a phenomenon of infinite CIFSs with the OSC which cannot hold in any finite CIFSs with the OSC.
In our previous papers [3] and [4], we considered a family of CIFSs of generalized complex continued fractions in order to present other examples of infinite CIFSs. In particular, we estimated the Hausdorff dimension of the limit set of each CIFS of the generalized complex continued fractions, and showed the non-positivity of the Hausdorff measure and the positivity of the packing measure of the limit set with respect to the Hausdorff dimension of the limit set. Note that the family of the CIFSs introduced in the papers [3] and [4] has uncountably many elements. However, there are unsolved problems in [3] and [4]: the estimation of the packing dimension and the proper-dimensional packing measure of the limit set. Therefore, the aim of this paper is to estimate the packing dimension and the proper-dimensional packing measure of each limit set in the family of the CIFSs of the generalized complex continued fractions, in order to improve the previous works [3] and [4].
@@
By the general theory of finite CIFSs with the OSC, the proper-dimensional Hausdorff measure and the proper-dimensional packing measure of the limit set are positive and finite (see [1], [2] and [7]). However, Corollary 1.5 indicates that for each $S_{\tau}$ in the family of CIFSs of generalized complex continued fractions which consists of uncountably many elements, the packing dimension of the limit set equals the Hausdorff dimension of the limit set, the proper-dimensional Hausdorff measure of the limit set is zero, and the proper-dimensional packing measure of the limit set is positive and finite. This is the phenomenon which Mauldin and Urbański first discovered in [7] and cannot hold in the finite CIFSs under the OSC.
@@
$m_{\tau}(B(x,r))$ $\geq m_{\tau}(B(0,cr))\geq m_{\tau}\left(\bigcup_{a\in I_{\tau}(r)}\phi_{a}(X)\right)\gtrsim\sum_{a\in I_{\tau}(r)}\ |a|^{-2h_{\tau}}$
#$\gtrsim\#(I_{\tau}(r))\cdot r^{2h_{\tau}}\gtrsim r^{2h_{\tau}-2}>r^{h_{\tau}},$
@@ 3 CIFSs of generalized complex continued fractions
$|\phi_{a}(z)|\cdot|a|=\frac{|a|}{|a+z|}=\sqrt{\frac{|a|^{2}}{|a+z|^{2}}}=\sqrt{\frac{|\phi_{a}^{\prime}(z)|}{|\phi_{a}^{\prime}(0)|}}\leq\sqrt{C}\leq K_{0}\leq K$